# Machine Unlearning Meets Adversarial Robustness via Constrained Interventions on LLMs

### v1

@@ 1 Introduction
Ensuring the safety of generative model outputs is a critical requirement for their deployment in real-world and safety-sensitive applications such as online content moderation and confidential data processing. To mitigate the risks of generating toxic or harmful content, prior work has explored strategies such as fine-tuning krause2021gedi ; dengresidual , prompt engineering luo2023prompt ; xie2023defending , and model augmentation dai2022knowledge ; meng2022locating . However, fine-tuning and model augmentation are computationally expensive, while prompt-based defenses are often brittle and sensitive to changes in model, context, and adversarial manipulation.
@@
Conversely, the task of unlearning sensitive or confidential content has mainly been addressed through partial retraining in teacher–student frameworks chen2023unlearn or iterative fine-tuning yao2024large ; tian2024forget , both of which incur substantial computational cost. Thus, there remains a gap for methods that jointly address robustness and unlearning while remaining computationally efficient. In this work, we bridge this gap by introducing a simple yet effective approach that simultaneously tackles adversarial robustness and unlearning. Specifically, our contributions can be summarized as follows:
@@ 4.1 Adversarial Robustness
Table 3: Attack success rate across models and formulations (One layer)
@@
Llama-3.1 8B 11.0 0.0 7.245 2.03 1.98 0.8
Mistral 7B v0.2 30.0 5.88 18.9 23.1 19.8 28.5
Gemma 2B-IT 22.0 2.508 8.225 12.98 12.76 19.58

### v2

@@ 1 Introduction
Ensuring the safety of generative model outputs is a critical requirement for their deployment in real-world and safety-sensitive applications such as online content moderation and confidential data processing. To mitigate the risks of generating toxic or harmful content, prior work has explored strategies such as fine-tuning krause2021gedi ; dengresidual , prompt engineering luo2023prompt ; xie2023defending , uncertainty quantification dakhmouche2025can and model augmentation dai2022knowledge ; meng2022locating . However, fine-tuning and model augmentation are computationally expensive, while prompt-based defenses are often brittle and sensitive to changes in model, context, and adversarial manipulation.
@@
Conversely, the task of unlearning sensitive or confidential content has mainly been addressed through partial retraining in teacher–student frameworks chen2023unlearn or iterative fine-tuning yao2024large ; tian2024forget , both of which incur substantial computational cost. Thus, there remains a gap for methods that jointly address robustness and unlearning while remaining computationally efficient. In this work, inspired by principled robustness approaches for regression dakhmouche2025noise; dakhmouche2025network; dakhmoucherobust, we bridge this gap by introducing a simple yet effective approach that simultaneously tackles adversarial robustness and unlearning. Specifically, our contributions can be summarized as follows:
@@ 4.1 Adversarial Robustness
Table 3: Attack Success Rate with different defenses - Custom Dataset- Lower better
@@
Llama-3.1 8B 60.50 5.49 30.17 3.00 2.00 18.00
Mistral 7B v0.2 90.00 36.47 61.00 77.00 66.00 95.00
Gemma 2B-IT 85.30 30.50 56.20 59.00 58.00 89.00